Write the following equations in the form y = mx + b and determine the number of solutions for each given system of equations.
(1) 3x + 9y = 36. (2) 2/3 + 2y = 8.
Number of solutions.
Convert each of the following equations to the form y = mx + b and verify whether the ordered pair (3, -2) is a solution - pair of the given system.
(1) 2x - 7y = 20. (2) x/4 + y/3 -2 = 0.
Is the ordered pair (3, -2) a solution - pair of this system?

Answer by Boreal(15235)

(1) 3x + 9y = 36. (2) 2/3 + 2y = 8.
(1) is 9y=-3x+36, and y=(-x/3)+4
(2) is 2y=22/3, y=11/3
One solution; 11/3=-x/3 +4; -x/3=-1/3; x=1 and y=11/3
----------------------------------
(1) 2x - 7y = 20. (2) x/4 + y/3 -2 = 0.
(1) is -7y=-2x+20; y=(2x/7)-(20/7), if x=3, then y=(6/7)-(20/7) and y=-2. That works.
(2) is y/3=-x/4 +2; y=(-3x/4)+6. If x=3, then y=15/4. that doesn't work.